OpenStudy (anonymous):

Factor the special product 64w^2-112w+49

OpenStudy (hoblos):

64w^2-112w+49 = (8w)^2 - 112w + 7^2 use the formula a^2 - 2ab + b^2 = (a-b)^2 then 64w^2-112w+49 = (8w-7)^2